# Smart Expert System: Large Language Models as Text Classifiers

### v1

@@ Abstract
Text classification is a fundamental task in Natural Language Processing (NLP), and the advent of Large Language Models (LLMs) has revolutionized the field. This paper introduces the Smart Expert System, a novel approach that leverages LLMs as text classifiers. The system simplifies the traditional text classification workflow, eliminating the need for extensive preprocessing and domain expertise. The performance of several LLMs, machine learning (ML) algorithms, and neural network (NN) based structures is evaluated on four datasets. Results demonstrate that certain LLMs surpass traditional methods in sentiment analysis, spam SMS detection and multi-label classification. Furthermore, it is shown that the system’s performance can be further enhanced through few-shot or fine-tuning strategies, making the fine-tuned model the top performer across all datasets. Source code and datasets are available in this GitHub repository: https://github.com/yeyimilk/llm-zero-shot-classifiers.
@@
Large Language Models, Text Classification, Natural Language Processing, Smart Expert Systems, Few-Shot Learning, Fine-Tuning, Chat GPT-4, Llama3.
@@
I Introduction
Text classification is a key task in NLP, widely used in areas like sentiment analysis Liu (2022), topic labeling Chen et al. (2020), question answering Minaee et al. (2021), and dialog act classification Qin et al. (2020). Traditional ML methods, such as logistic regression, Naive Bayes, and support vector machines, are common but need a lot of labeled data and can only handle known classes Sarker (2021); Wang et al. (2019). Deep learning (DL) methods, including deep neural networks, recurrent neural networks, and convolutional neural networks, have surpassed traditional ML algorithms by capturing complex data relationships LeCun et al. (2015); Sutskever et al. (2011). However, DL methods also require large labeled datasets and face challenges in adapting to new tasks without significant retraining.
LLMs, especially those based on the Transformer architecture like PaLM Chowdhery et al. (2023), LLaMA Touvron et al. (2023), GPT Radford et al. (2018), and BERT Devlin et al. (2018), have revolutionized NLP. With hundreds of billions to trillions of parameters, these models are pre-trained on massive amounts of text data. Compared to traditional DL methods, LLMs are much larger and exhibit stronger language understanding and generation capabilities. They learn rich language representations through extensive pretraining, enabling them to adapt quickly to various downstream tasks. For instance, Přibáň et al. (2024) used LLMs for cross-lingual sentiment analysis. Caruccio et al. (2024) explored using LLMs for intelligent diagnoses. DePalma et al. (2024) examined LLMs’ code refactoring capabilities. Additionally, Boubker (2024) and Oh et al. (2024) evaluated and enhanced student learning performance using LLMs.
Although LLMs have been widely applied and researched for various tasks, their potential for handling text classification tasks remains to be further explored. Recent studies indicate that while LLMs perform excellently in syntactic analysis, there are still numerous challenges in achieving efficient and accurate text classification in practical applications. The studies by Bețianu et al. (2024) and Møller et al. (2024) demonstrate the application of LLMs in domain adaptation and synthetic data augmentation, revealing both the strengths and limitations of these models in handling rare categories or domain shifts. However, the development of LLMs offers new research opportunities to simplify and optimize the text classification process.
Traditional ML and NN approaches to text classification involve a multi-stage pipeline that can be both complex and resource-intensive. These stages typically include feature extraction, dimensionality reduction, classifier selection, and model evaluation Kowsari et al. (2019). Figure 1 illustrates this conventional workflow, which starts with data collection followed by laborious preprocessing steps such as tokenization, punctuation removal, stop word filtering, stemming, and lemmatization. Subsequently, feature extraction techniques like bag-of-words, word embeddings, or TF-IDF are applied before training a classifier to produce the final classification model.
However, this traditional approach presents several challenges. Firstly, it requires significant domain expertise to effectively preprocess the data and engineer features that capture the relevant information for the classification task. Secondly, each step in the process must be carefully tuned to optimize performance—a time-consuming endeavor that often involves trial-and-error experimentation. Lastly, these methods may not generalize well across different datasets or languages without substantial reconfiguration.
In contrast, LLMs offer a much easier way to do text classification, as depicted in Figure 2, that it just includes three main blocks: data collection, feeding data directly into LLMs, and obtaining classification results from LLMs. This simplified flow eliminates the need for explicit feature extraction or dimensionality reduction since LLMs inherently encode rich linguistic features through their deep contextual representations. Moreover, because they have been pre-trained on diverse datasets, they require minimal—if any—additional training specific to a given task or domain.
In light of these considerations, we propose a novel approach that capitalizes on the strengths of LLMs for text classification tasks. Our approach is encapsulated in the Smart Expert System, which utilizes LLMs as Text Classifiers, as depicted in Figure 3. This system is designed to leverage the advanced capabilities of LLMs, offering a more streamlined and efficient method for classifying text without the extensive preprocessing and feature engineering required by traditional ML/NN methods. Introducing our Smart Expert System represents a significant shift towards more intelligent and adaptable NLP tools. By integrating LLMs into the core of text classification workflows, we aim to reduce complexity and enhance performance across various domains and applications.
@@ I Introduction
  • This paper proposes a text classification system using LLMs to replace traditional text classifiers. This system simplifies the conventional text classification process, lowering technical barriers and eliminating the need for domain experts to perform complex preprocessing and algorithm design. This approach is crucial for rapid deployment and scalable applications, especially for small businesses that do not have deep expertise in ML or DL.
@@
  • We compare the performance of LLMs with traditional ML and NN models across multiple datasets. The results show that LLMs, after learning from a small number of samples or fine-tuning, outperform in various text classification tasks. This finding confirms the versatility and efficiency of LLMs.
Figure 3: Framework of our Expert System. The steps of the framework can be included as (1) collect data from the data source to establish the domain database; (2) send domain-specific data to the pre-trained LLM model, like GPT-4, Llama-3 and so on; (3) using a few domain-specific data to do fine-tuning or few-shot learning; (4) apply the fine-tuning or few-shot learning to the pre-trained LLM model; (5) (optional) utilize domain knowledge expert to set up the prompts to elevate LLM performance; (6) apply prompts in the pre-trained model; (7) evaluate the whole expert system’s performance; (8) non-expert users query tasks through user interface to the expert system; (Tasks may include classification, sentiment analysis, prediction, recommendation and so on. In this paper, we take the classification and sentiment analysis as examples.) (9) LLM API interacts with User interface and the pre-trained LLM model, advising on the user interface.
@@ II-A Traditional Text Classification Approaches
Text classification has evolved through several methodologies, each with unique advantages and limitations. Early rule-based methods like decision trees and expert systems, such as C4.5 Quinlan (2014) and MYCIN Shortliffe (2012), were straightforward but prone to overfitting and inflexibility with new or noisy data. In contrast, probability-based models like Naive Bayes Xu (2018) and Hidden Markov Models Rabiner (1989) offered better generalization for tasks like spam detection and speech recognition, handling sequential and complex data effectively. Geometry-based approaches, including SVMs Joachims (2002) and dimensionality reduction techniques like PCA Wold et al. (1987), excelled in managing high-dimensional spaces but sometimes struggled with large datasets and computational efficiency. Finally, statistic-based methods like KNN Guo et al. (2003) and Logistic Regression Genkin et al. (2007) provided insights through nearest neighbor analysis and probability estimation, although they were sensitive to data preprocessing and could falter with nonlinear data.
@@ II-B Deep Learning Approaches
Deep learning has become a key technology in text classification, capable of handling complex language features. CNN text classification models Kim (2014) capture local textual features through convolutional layers. LSTM Sutskever et al. (2014) and GRU Chung et al. (2014), as optimized versions of RNNs, are particularly effective in addressing long-distance dependencies in text. Zhou et al. (2024b) proposes an optical character recognition and classification method for cigarette laser code recognition, using a convolutional recurrent neural network to extract image features and utilizing BiLSTM for text classification. Transformer models, like BERT Devlin et al. (2018), achieve remarkable results in various NLP tasks by utilizing self-attention mechanisms. Specifically, the BERT model demonstrates powerful capabilities in text classification tasks. Raza et al. (2024) proposes the text recognition framework Nbias for detecting and eliminating biases, including data, corpus construction, model development, and evaluation layers. The dataset is collected from various fields, and a transformer-based token classification model is applied. Zhou et al. (2024a) proposes a semi-supervised generative adversarial learning method that improves the model’s classification performance with limited annotated data through generative adversarial networks. Jamshidi et al. (2024) introduces a hybrid model that combines BERT, LSTM, and Decision Templates (DT) for IMDB and Drug Review classification. However, these methods typically require substantial data for training, often necessitating extensive datasets to achieve optimal performance. This reliance on large training sets can pose challenges, especially when collecting or labeling data is difficult or impractical.
@@ II-C LLM Approaches
LLMs represent a significant advancement in the field of text classification, building on the DL foundations that have revolutionized NLP. These models, which include notable examples such as GPT Brown et al. (2020), T5 Raffel et al. (2020), RWKV Peng et al. (2023), Mamba Gu & Dao (2023), Gemini Team et al. (2023), PaLM Chowdhery et al. (2023), Llama Touvron et al. (2023), and Claude Anthropic (2024), leverage massive amounts of data and extensive training regimes to understand and generate human-like text. Their ability to capture nuanced language patterns and context makes them highly effective for text classification tasks across various domains.
Recent studies have begun to explore the practical applications of LLMs in specialized fields. For instance, Chae & Davidson (2023) study the application of LLMs in sociological text classification, demonstrating their potential in social science research. Loukas et al. (2023) examine the performance and cost trade-offs when employing LLMs for text classification, focusing on financial intent detection datasets. Another study by Wei et al. (2023) investigates the effect of fine-tuning LLMs on text classification tasks within legal document review, highlighting how domain-specific adjustments can enhance model performance. Furthermore, the research identified as Cloutier & Japkowicz (2023) refines LLM performance on multi-class imbalanced text classification tasks through oversampling techniques, addressing one of the common challenges in ML.
Despite these advancements, there remains a gap between state-of-the-art methods and their accessibility to users without deep technical expertise or resources for extensive model training and tuning. The proposed smart expert system aims to bridge this gap by utilizing pre-trained LLMs as out-of-the-box classifiers that require minimal additional processing or domain-specific adaptation. This approach simplifies the deployment process and democratizes access to advanced NLP tools, enabling a wider range of users to benefit from cutting-edge technology without prohibitive costs or steep learning curves.
In summary, while traditional ML and NN methods have necessitated laborious preprocessing steps and dataset-specific model training, pre-trained LLMs offer an elegant solution that bypasses these requirements without compromising accuracy or effectiveness. By harnessing their inherent understanding of language nuances and context sensitivity, pre-trained LLMs can be directly applied to diverse text classification scenarios with remarkable efficiency—ushering in a new era where sophisticated NLP capabilities are readily available to all who seek them.
@@ III Methodology
III-A Smart Expert System
Our proposed system integrates LLMs to refine the traditional text classification system, as illustrated in Figure 3. The framework of our expert system presents a comprehensive strategy that capitalizes on the strengths of LLMs while mitigating their traditional limitations.
@@ III-A Smart Expert System
Furthermore, the involvement of a domain knowledge expert is crucial but optional. They configure the system by establishing customized prompts that direct the LLM toward generating pertinent and contextually aware responses. This human-in-the-loop methodology guarantees that the system adheres to specific domain requirements and can adeptly manage intricate query scenarios.
@@
By amalgamating these components, our Smart Expert System simplifies and elevates text classification processes in terms of adaptability and precision. It significantly diminishes reliance on domain experts for complex preprocessing or algorithmic design tasks—thereby democratizing access to cutting-edge NLP technologies across various sectors such as e-commerce and social media analytics.
@@ III-B2 F1 Score
First, we define precision and recall:
$\text{Precision}=\frac{TP}{TP+FP}$ (2)
$\text{Recall}=\frac{TP}{TP+FN}$ (3)
Then, the F1 score is given by:
$F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (4)
@@ IV Dataset
The empirical evaluation of zero-shot text classification capabilities of LLMs in this study is conducted across four datasets, each chosen for its unique characteristics and relevance to specific domains. The datasets are as follows:
@@ IV-A COVID-19-related Tweets Dataset
The first dataset consists of tweets related to the COVID-19 pandemic, curated by Preda (2020). This dataset is particularly valuable for sentiment analysis due to the real-time nature of Twitter data and the global impact of the pandemic, which has elicited a wide range of sentiments from users worldwide. As shown in Table I, it comprises a total of 41,156 training instances and 3,798 test instances, categorized into negative, neutral, and positive sentiments.
@@ IV-B Economic Texts Dataset
The second dataset includes economic texts compiled by Malo et al. (2014), designed for sentiment analysis within the financial domain. Economic texts present different linguistic features than social media content, often containing specialized vocabulary and complex expressions. The dataset contains 3,876 training samples and 969 test samples distributed across negative, neutral, and positive classes, as detailed in Table II. This dataset includes 5 levels of sentiment, which were merged into 3 levels in this study.
@@ IV-C E-commerce Texts Dataset
For multi-class classification tasks beyond binary or ternary sentiment analysis, we utilize an e-commerce texts dataset provided by Gautam (2019). This dataset encompasses product reviews from four categories: Household items, Books, Clothing & Accessories (C&A), and Electronics. It offers a practical application scenario where accurate classification can directly influence recommendation systems and customer satisfaction. The training set includes 40,338 instances, while the test set contains 10,085 instances, as outlined in Table III.
@@ IV-D SMS Spam Collection Dataset
Lastly, we incorporate an SMS Spam Collection dataset assembled by Almeida & Hidalgo (2012) to evaluate spam detection performance. Spam detection is critical for maintaining communication integrity and user trust in messaging platforms. This binary classification task involves distinguishing between normal messages and spam with a total of 4,457 training messages and 1,115 test messages presented in Table IV.
Each dataset was selected not only for its relevance to current societal issues but also because it collectively provides a comprehensive assessment environment that spans various text classification challenges faced by LLMs today. These include varying lengths of text inputs (from short tweets to longer reviews), domain-specific language usage (as seen in economic texts), diverse sentiment expressions (ranging from public health concerns to consumer products), and practical applications such as spam filtering. The diversity among these datasets will enable us to evaluate zero-shot learning capabilities across different contexts using LLMs robustly.
@@ V-A Experiment Setup
  • ZSL Models: We explore zero-shot learning capabilities using transformer-based models, including BART (facebook/bart-large-mnli) and DeBERTa (microsoft/deberta-large-mnli).
@@
Conversely, unprocessed raw text from the testing dataset was used to fully leverage their natural language understanding abilities for zero-shot learning models and LLMs. It is important to note that this testing dataset remains consistent across all model types to provide a fair comparison platform.
@@
By adopting these methodological choices—standardized input processing for traditional ML algorithms/NN architectures and careful instance selection respecting original distributions—we aim to provide an equitable experimental environment that allows us to assess each model’s inherent zero-shot text classification capabilities under comparable conditions.
@@ V-B Experimental Results
The following tables present the experimental results for all the models. Notably, when employing few-shot strategies or fine-tuning, they are indicated by ”(S)” in the table or denoted by ”(F).” The value within the parenthesis shows the difference between the base LLM model result and the model employed with the S or F strategy.
@@
Traditional algorithms, such as MNB, LR, RF, DT, and KNN, show poor accuracy and F1 scores, with RF performing marginally better than others in its class. In contrast, NN-based models demonstrate superior performance, with GRU leading in both ACC and F1 metrics. Among LLMs, before fine-tuning, GPT-3.5 exhibits the highest ACC and F1 scores, outperforming other LLMs, including GPT-4, while the performance is below NN methods. However, once the fine-tuning method was employed, the Qwen-7B(F) outperformed all the other models, including GRU’s best model. The best accuracy increased from 0.6913, performed by GRU, to 0.8388, and the F1 score from 0.63332, performed by RNN, to 0.8433.

### v2

# Adaptable and Reliable Text Classification using Large Language Models

@@ Abstract
Text classification is fundamental in Natural Language Processing (NLP), and the advent of Large Language Models (LLMs) has revolutionized the field. This paper introduces an adaptable and reliable text classification paradigm, which leverages LLMs as the core component to address text classification tasks. Our system simplifies the traditional text classification workflows, reducing the need for extensive preprocessing and domain-specific expertise to deliver adaptable and reliable text classification results. We evaluated the performance of several LLMs, machine learning algorithms, and neural network-based architectures on four diverse datasets. Results demonstrate that certain LLMs surpass traditional methods in sentiment analysis, spam SMS detection, and multi-label classification. Furthermore, it is shown that the system’s performance can be further enhanced through few-shot or fine-tuning strategies, making the fine-tuned model the top performer across all datasets. Source code and datasets are available in this GitHub repository: https://github.com/yeyimilk/llm-zero-shot-classifiers.
@@
Large Language Models, Text Classification, Natural Language Processing, Adaptive Learning, Fine-Tuning, Chat GPT-4, Llama3.
I Introduction
Text classification is a core task in natural language processing (NLP), with applications ranging from sentiment analysis to question answering [1, 2, 3]. Traditional machine learning (ML) methods, such as logistic regression and Naive Bayes [4, 5], have been widely employed. However, these approaches often require extensive labeled datasets and are limited in adapting to unseen data or emerging categories, thus posing challenges in dynamic real-world environments.
The emergence of large language models (LLMs) based on Transformer architectures, such as PaLM [6], LLaMA [7], and GPT [8], has transformed the landscape of text classification. Unlike traditional approaches, as shown in Figure 1, which require complex, multi-step pipelines for data preprocessing and feature extraction, LLMs leverage their extensive pre-training to handle these tasks internally. In contrast to the more labor-intensive traditional method, this shift reduces the need for manual intervention and allows the models to generalize more effectively across various domains. As illustrated in Figure 2, the LLM-based approach condenses the workflow into three main stages: data collection, feeding data directly into the LLM, and receiving classification outputs.
@@ I Introduction
Figure 3: Framework of our adaptable and reliable text classification system. The steps of the framework can be included as (1) collect data from the data source to establish the domain database; (2) send domain-specific data to the pre-trained LLM model, like GPT-4, Llama-3 and so on; (3) using a few domain-specific data to do fine-tuning or few-shot learning; (4) apply the fine-tuning or few-shot learning to the pre-trained LLM model; (5) (optional) utilize domain knowledge to set up the prompts to elevate LLM performance; (6) apply prompts in the pre-trained model; (7) evaluate the whole system’s performance; (8) non-expert users query tasks through user interface to the system; (Tasks may include classification, sentiment analysis, prediction, recommendation and so on. In this paper, we take the multi-class classification and sentiment analysis as examples.) (9) LLM API interacts with User interface and the pre-trained LLM model, advising on the user interface.
Despite these advancements, deploying LLMs in real-world text classification tasks still presents particular challenges. For instance, LLMs must maintain high reliability across diverse and unpredictable environments, ensuring robustness even when dealing with domain shifts or rare categories [9, 10]. Additionally, the responsibility of these models is increasingly essential, as fairness, transparency, and ethical considerations come to the forefront when implementing LLMs in decision-making systems [11].
In response to these challenges, we propose a novel text classification framework that harnesses the strengths of LLMs while addressing critical aspects of adaptability. Our framework, illustrated in Figure 3, integrates LLMs at the core of the classification workflow, significantly simplifying the process. It allows non-expert users to access high-performing classification systems with minimal effort, reducing the extensive preprocessing and feature engineering traditionally required.
@@
  • This paper proposes a text classification system using LLMs to replace traditional text classifiers. This system simplifies the conventional text classification process, lowering technical barriers and eliminating the need for domain experts to perform complex preprocessing and algorithm design. This approach is crucial for rapid deployment and scalable applications, especially for small businesses needing deep ML or DL expertise.
@@
  • We compare the performance of LLMs with traditional ML and NN models across multiple datasets. After learning from a few samples or fine-tuning, the results show that LLMs outperform in various text classification tasks. This finding confirms the versatility and efficiency of LLMs.
@@ II-A Traditional Text Classification Approaches
Text classification has evolved through various machine learning (ML) methods, each with its strengths and limitations. Early rule-based approaches, such as decision trees like C4.5 [12], were simple but prone to overfitting and lacked flexibility. Probability-based models, such as Multinomial Naive Bayes (MNB) [13] and Hidden Markov Models [14], improved generalization, particularly in tasks like spam detection and speech recognition. Geometry-based methods, including support vector machine (SVM) [15], handled high-dimensional data but struggled with large datasets. Finally, statistical methods like K-nearest neighbors (KNN) [16] and Logistic Regression (LG) [17] provided effective solutions but required extensive preprocessing and often faltered with nonlinear data.
@@ II-B Deep Learning Approaches
Deep learning (DL) has become a key technology in text classification, capable of handling complex language features. Convolutional neural networks (CNNs) text classification models [18] capture local textual features through convolutional layers. LSTM [19] and GRU [20], as optimized versions of RNNs, are particularly effective in addressing long-distance dependencies in text. [21] proposes an optical character recognition and classification method for cigarette laser code recognition, using a convolutional recurrent neural network to extract image features and utilizing BiLSTM for text classification. Transformer models, like BERT [22], achieve remarkable results in various NLP tasks by utilizing self-attention mechanisms. Specifically, the BERT model demonstrates powerful capabilities in text classification tasks. [23] proposes the text recognition framework Nbias for detecting and eliminating biases, including data, corpus construction, model development, and evaluation layers. The dataset is collected from various fields, and a transformer-based token classification model is applied. [24] proposes a semi-supervised generative adversarial learning method that improves the model’s classification performance with limited annotated data through generative adversarial networks. [25] introduces a hybrid model that combines BERT, LSTM, and Decision Templates (DT) for IMDB and Drug Review classification. However, these methods typically require substantial data for training, often necessitating extensive datasets to achieve optimal performance. This reliance on large training sets can pose challenges, especially when collecting or labeling data is difficult or impractical.
@@ II-C LLM Approaches
LLMs represent a significant advancement in the field of text classification, building on the DL foundations that have revolutionized NLP. These models, which include notable examples such as GPT [26], T5 [27], RWKV [28], Mamba [29], Gemini [30], PaLM [6], Llama [7], and Claude [31], leverage massive amounts of data and extensive training regimes to understand and generate human-like text. Their ability to capture nuanced language patterns and context makes them highly effective for text classification tasks across various domains.
Recent studies have begun to explore the practical applications of LLMs in specialized fields. For instance, [32] study the application of LLMs in sociological text classification, demonstrating their potential in social science research. [33] examine the performance and cost trade-offs when employing LLMs for text classification, focusing on financial intent detection datasets. Another study by [34] investigates the effect of fine-tuning LLMs on text classification tasks within legal document review, highlighting how domain-specific adjustments can enhance model performance. Furthermore, the research identified as [35] refines LLM performance on multi-class imbalanced text classification tasks through oversampling techniques, addressing one of the common challenges in ML.
Despite their strengths, there remains a gap in making LLMs accessible to users without deep technical expertise. Our system addresses this by leveraging pre-trained LLMs as out-of-the-box classifiers that require minimal adaptation. This system democratizes access to advanced NLP tools, offering scalable solutions for diverse applications without the steep learning curve typically associated with LLM deployment.
@@ III Methodology
III-A Adaptable and Reliable System
Our proposed system integrates LLMs to refine the traditional text classification system, as illustrated in Figure 3 based on our previous work [36]. The framework of our system presents a comprehensive strategy that capitalizes on the strengths of LLMs while mitigating their traditional limitations.
@@ III-A Adaptable and Reliable System
Furthermore, the involvement of domain knowledge is crucial but optional. They configure the system by establishing customized prompts that direct the LLM toward generating pertinent and contextually aware responses. This human-in-the-loop methodology guarantees that the system adheres to specific domain requirements and can adeptly manage intricate query scenarios.
@@
By amalgamating these components, our system simplifies and elevates text classification processes in terms of adaptability and precision. It significantly diminishes reliance on domain knowledge for complex preprocessing or algorithmic design tasks—thereby democratizing access to cutting-edge NLP technologies across various sectors such as e-commerce and social media analytics.
@@ III-B2 F1 Score
$F1=\frac{\text{2TP}}{\text{2TP}+\text{FP}+\text{FN}}$ (2)
@@ IV Dataset
Four datasets include varying lengths of text inputs (from short tweets to longer reviews), domain-specific language usage (as seen in economic texts), diverse sentiment expressions (ranging from public health concerns to consumer products), and practical applications such as spam filtering, were employed to evaluate the LLMs’ adaptability and reliability in handling text classification tasks.
@@ IV-A COVID-19-related Tweets Dataset
The first dataset consists of tweets related to the COVID-19 pandemic, curated by [37]. As shown in Table I, it comprises a total of 41,156 training instances and 3,798 test instances, categorized into negative, neutral, and positive sentiments.
@@ IV-B Economic Texts Dataset
The second dataset includes economic texts compiled by [38], designed for sentiment analysis within the financial domain. The dataset contains 3,876 training samples and 969 test samples distributed across negative, neutral, and positive classes, as detailed in Table II. This dataset includes 5 levels of sentiment, which were merged into 3 levels in this study.
@@ IV-C E-commerce Texts Dataset
For multi-class classification tasks beyond binary or ternary sentiment analysis, we utilize an e-commerce texts dataset provided by [39]. The training set includes 40,338 instances, while the test set contains 10,085 instances, as outlined in Table III.
@@ IV-D SMS Spam Collection Dataset
Lastly, we incorporate an SMS Spam Collection dataset assembled by [40] to evaluate spam detection performance. This binary classification task involves distinguishing between normal messages and spam with a total of 4,457 training messages and 1,115 test messages presented in Table IV.
@@ V-A Experiment Setup
  • Zero-shot Learning (ZSL) Models: We explore zero-shot learning capabilities using transformer-based models, including BART (facebook/bart-large-mnli) and DeBERTa (microsoft/deberta-large-mnli).
@@
Conversely, unprocessed raw text from the testing dataset was used to fully leverage their natural language understanding abilities for zero-shot learning models and LLMs. It is important to note that this testing dataset remains consistent across all model types to provide a fair comparison.
@@ V-B Experimental Results
Table V, VI, VII and VIII present the experimental results for all the models. Notably, when employing few-shot strategies or fine-tuning, they are indicated by “(S)” and “(F),” respectively.
@@
Traditional algorithms show poor accuracy and F1 scores. In contrast, NN-based models demonstrate superior performance, with GRU leading in both ACC and F1 metrics. Among LLMs, before fine-tuning, GPT-3.5 exhibits the highest ACC and F1 scores, outperforming other LLMs, including GPT-4, while the performance is below NN methods. However, once the fine-tuning method was employed, the Qwen-7B(F) outperformed all the other models, including GRU’s best model. The best accuracy increased from 0.6913, performed by GRU, to 0.8388, and the F1 score from 0.63332, performed by RNN, to 0.8433.